• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 14:33:18 on September24,2004

    Entry number 131865

    This entry is a followup to: 131690

    keyword=BCM cabling fixed

    Arun & Bob: The BCM signals were fixed. The d3 on L-arm was fixed
    when we re-plugged in the cables upstairs. The cable does not appear
    to be bad, I suppose someone had not plugged it in all the way.
    The u10 on R-arm was bad because of the RG213 cable from the Hall A
    floor to the back of the detector stack. Cable #7 is bad. We sent
    the signal a different route.